# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ Abstract.
This is a continuation of [16] studying the discrete conformal structures on surfaces with boundary, in which we gave a classification of the discrete conformal structures on surfaces with boundary. In this paper, we prove the rigidity and existence of these discrete conformal structures on surfaces with boundary.
@@ 1. Introduction
1.1. Basic definition and notations
@@
Let $\widetilde{\mathcal{T}}={(V,\widetilde{E},\widetilde{F})}$ be a triangulation of a closed surface $\widetilde{\Sigma}$, where $V,\widetilde{E},\widetilde{F}$ represent the sets of vertices, edges and faces respectively and $V$ is a finite subset of $\widetilde{\Sigma}$ with $|V|=N$. Denote $N(V)$ as a small open regular disjoint neighborhood of the union of all vertices $V$. Then $\Sigma=\widetilde{\Sigma}-N(V)$ is a compact surface with $N$ boundary components. The intersection $\mathcal{T}=\widetilde{\mathcal{T}}\cap\Sigma$ is called an ideal triangulation of the surface $\Sigma$. The intersections $E=\widetilde{E}\cap\Sigma$ and $F=\widetilde{F}\cap\Sigma$ are the sets of the ideal edges and ideal faces of $\Sigma$ in the ideal triangulation $\mathcal{T}$ respectively. The intersection of an ideal face and the boundary $\partial\Sigma$ is called a boundary arc. For simplicity, we use $B=\{1,2,...,N\}$ to denote the boundary components and label them as $i\in B$. We use $E_{+}$ to denote the set of oriented edges. Denote $\{ij\}\in E$ as the ideal edge between two adjacent boundary components $i,j\in B$. Label the ordered ideal edge as $(i,j)$. Denote $\{ijk\}\in F$ as the ideal face adjacent to the boundary components $i,j,k\in B$. In fact, the ideal face $\{ijk\}$ is a hexagon, which corresponds to the triangle $v_{i}v_{j}v_{k}\in\widetilde{F}$ in $\widetilde{\mathcal{T}}$. The sets of real valued functions on $B,E$ and $E_{+}$ are denoted by $\mathbb{R}^{N}$, $\mathbb{R}^{E}$ and $\mathbb{R}^{E_{+}}$ respectively.
@@
An edge length function associated to $\mathcal{T}$ is a vector $l\in\mathbb{R}^{E}_{>0}$ assigning each ideal edge $\{ij\}\in E$ a positive number $l_{ij}=l_{ji}$. For any three positive numbers, there exists a unique right-angled hyperbolic hexagon (up to isometry) with the lengths of three non-pairwise adjacent edges given by these three positive numbers [7]. Hence, for any ideal face $\{ijk\}\in F$, there exists a unique right-angled hyperbolic hexagon whose three non-pairwise adjacent edges have lengths $l_{ij},l_{ik},l_{jk}$. Gluing all such right-angled hyperbolic hexagons isomorphically along the ideal edges in pairs, one can construct a hyperbolic surface with totally geodesic boundary from the ideal triangulation $\mathcal{T}$. Conversely, any ideally triangulated hyperbolic surface $(\Sigma,\mathcal{T})$ with totally geodesic boundary produces a function $l\in\mathbb{R}^{E}_{>0}$ with $l_{ij}$ given by the length of the shortest geodesic connecting the boundary components $i,j\in B$. The edge length function $l\in\mathbb{R}^{E}_{>0}$ is called a discrete hyperbolic metric on $(\Sigma,\mathcal{T})$. The length $K_{i}$ of the boundary component $i\in B$ is called the generalized combinatorial curvature of the discrete hyperbolic metric $l$ at $i$. Namely, the generalized combinatorial curvature $K_{i}$ is defined to be
$K_{i}=\sum_{\{ijk\}\in F}\theta^{jk}_{i},$
@@
where the summation is taken over all the right-angled hyperbolic hexagons adjacent to $i$ and the generalized angle $\theta^{jk}_{i}$ is the length of the boundary arc of the right-angled hyperbolic hexagon $\{ijk\}$ at $i$.
@@
Motivated by Glickenstein’s work in [2] and Glickenstein-Thomas’ work in [3] for triangulated closed surfaces, we introduced the following two definitions of partial edge length and discrete conformal structures on an ideally triangulated surface with boundary.
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. An assignment of partial edge lengths is a map $d\in\mathbb{R}^{E_{+}}$ such that $l_{ij}=d_{ij}+d_{ji}>0$ for every edge $\{ij\}\in E$ and
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. A discrete conformal structure $d=d(f)$ on $(\Sigma,\mathcal{T})$ is a smooth map, sending a function $f\in\mathbb{R}^{N}$ defined on boundary components $B$ to a partial edge length function $d\in\mathbb{R}^{E_{+}}$, such that
@@
In general, the partial edge lengths $d\in\mathbb{R}^{E_{+}}$ do not satisfy the symmetry condition, i.e., $d_{ij}\neq d_{ji}$. Denote $E_{ij}$ as the hyperbolic geodesic in the hyperbolic plane $\mathbb{H}^{2}$ extending the ideal edge $\{ij\}\in E$. The partial edge length $d_{ij}$ represents the signed distance from a point $c_{ij}$ in $E_{ij}$ to $i\in B$ and $d_{ji}$ represents the signed distance from $c_{ij}$ to $j\in B$. This point $c_{ij}$ is called the edge center of the ideal edge $\{ij\}$. Note that the signed distance $d_{ij}$ of $c_{ij}$ to $i$ is positive if $c_{ij}$ is on the same side of $i$ as $j$ along the hyperbolic geodesic $E_{ij}$, and negative otherwise. Please refer to Figure 1.
Figure 1. Poincaré dual to the ideal edge $\{ij\}\in E$ with the position determined by $d_{ij}$ and $d_{ji}$ in the Klein model (dotted). The red lines represent the boundary arcs of the ideal face $\{ijk\}$. The black triangle $v_{i}v_{j}v_{k}$ is the corresponding hyper-ideal hyperbolic triangle.
@@
We briefly explain the motivation of Definition 1.1 for the partial edge length on surfaces with boundary. One can refer to [16] for more details. The hyperbolic geodesic orthogonal to the hyperbolic geodesic $E_{ij}$ through $c_{ij}$ is called the edge perpendicular, denoted by $P_{ij}$. We can identify the perpendicular $P_{ij}$ with a unique 2-dimensional time-like subspace $\widetilde{P}_{ij}$ of the Lorentzian space $\mathbb{R}^{3}$ such that $\widetilde{P}_{ij}\cap\mathbb{H}^{2}=P_{ij}$. Since the 2-dimensional time-like subspaces $\widetilde{P}_{ij}$ and $\widetilde{P}_{jk}$ associated to $P_{ij}$ and $P_{jk}$ intersect in a 1-dimensional subspace of $\mathbb{R}^{3}$, then the intersection $\widetilde{P}_{ij}\cap\widetilde{P}_{jk}$ corresponds to a point $q$ in the Klein model of hyperbolic plane $\mathbb{H}^{2}$. We view this point $q$ as the intersection point of the perpendiculars $P_{ij}$ and $P_{jk}$. Note that the point $q$ may be not in $\mathbb{H}^{2}$, i.e., the perpendiculars $P_{ij},P_{jk}$ may not intersect within $\mathbb{H}^{2}$. Lemma 2.4 in [16] shows that the perpendiculars $P_{ij},P_{jk},P_{ki}$ of $\{ijk\}$ intersect at a common point $c_{ijk}$ (called face center) if and only if the condition (1) holds. Hence, the condition (1) in Definition 1.1 ensures the existence of a dual geometric structure on the Poincaré dual of an ideally triangulated surface with boundary.
Remark 1.5.
The geometric motivation for Definition 1.2 is that the vertex $v_{k}$ and face center $c_{ijk}$ of a hyper-ideal hyperbolic triangle should be in the 2-dimensional vector subspace of $\mathbb{R}^{3}$ including the vertex $\overline{v}_{k}$ of the new hyper-ideal hyperbolic triangle under the $f_{k}$-conformal variation. Please refer to [16, Remark 1.3] for more details.
The main result in [16] is the following theorem, which gives the explicit forms of the discrete conformal structures in Definition 1.2.
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$. There exist constant vectors $\alpha\in\mathbb{R}^{N}$, $\eta\in\mathbb{R}_{>0}^{E}$ and $\zeta\in\mathbb{R}^{E}$ satisfying $\eta_{ij}=\eta_{ji}$ and $\zeta_{ij}=\zeta_{ji}$, such that for any right-angled hyperbolic hexagon $\{ijk\}\in F$,
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$, then the discrete conformal structure $d=d(f)$ is
(4) $\coth d_{ij}=-\frac{\alpha_{i}e^{2f_{i}}}{\sinh l_{ij}}\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{i}e^{2f_{i}}}}+\frac{\eta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}>0$, then the discrete conformal structure $d=d(f)$ is
(6) $\coth d_{ij}=\frac{\sinh(f_{j}-f_{i}-C_{ij})}{\sinh l_{ij}}+\frac{\eta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(7) $\cosh l_{ij}=-\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}},$
where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$ and $C_{rs}+C_{sr}=0$ for any $\{r,s\}\subseteq\{i,j,k\}$;
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$, then the discrete conformal structure $d=d(f)$ is
(8) $\coth d_{ij}=\frac{\alpha_{i}e^{2f_{i}}}{\sinh l_{ij}}\sqrt{\frac{1+\alpha_{j}e^{2f_{j}}}{1+\alpha_{i}e^{2f_{i}}}}+\frac{\zeta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(9) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\zeta_{ij}e^{f_{i}+f_{j}};$
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}<0$, then the discrete conformal structure $d=d(f)$ is
$\coth d_{ij}=-\frac{\sinh(f_{j}-f_{i}-C_{ij})}{\sinh l_{ij}}+\frac{\zeta_{ij}e^{f_{i}+f_{j}}}{\sinh l_{ij}}$
with
(10) $\cosh l_{ij}=\cosh(f_{j}-f_{i}-C_{ij})+\zeta_{ij}e^{f_{i}+f_{j}},$
where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$ and $C_{rs}+C_{sr}=0$ for any $\{r,s\}\subseteq\{i,j,k\}$.
@@
As mentioned in [16, Remark 1.5], we can always assume $\alpha:B\rightarrow\{-1,0,1\}$ in the discrete conformal structures (A1) and (B1) and $C\equiv 0$ in the discrete conformal structures (A2) and (B2) for simplicity.
@@
According to [16, Remark 1.7], the discrete conformal structures (A1) and (A2) can exist alone on an ideally triangulated surface with boundary, but they can not exist simultaneously on the same ideally triangulated surface with boundary. Moreover, the discrete conformal structures (B1) and (B2) can not exist alone on an ideally triangulated surface with boundary. Therefore, we consider the mixed type of discrete conformal structures, i.e., the discrete hyperbolic metrics $l_{ij},l_{jk},l_{ki}$ of a right-angled hyperbolic hexagon $\{ijk\}\in F$ are induced by (7) and (10) or (5) and (9), such that the condition (1) holds. We call the former the mixed discrete conformal structure I and the latter the mixed discrete conformal structure II. For example, set $\{r,s,t\}=\{i,j,k\}$, if $d_{rs}<0$, then $d_{rt}<0$ by Lemma 2.17, and $d_{sr}>0,\ d_{tr}>0$ by the conditions $l_{rs}=d_{rs}+d_{sr}>0$ and $l_{rt}=d_{rt}+d_{tr}>0$ in Definition 1.1. We need $d_{st}>0,\ d_{ts}>0$ to make the condition (1) hold. Thus $l_{rs},l_{rt}$ are defined by (10) and $l_{st}$ is defined by (7) or $l_{rs},l_{rt}$ are defined by (9) and $l_{st}$ is defined by (5).
1.2. Rigidity and existence of discrete conformal structures
An important problem in discrete conformal geometry is to obtain the relationships between the discrete conformal structures and their combinatorial curvatures. The main result in this paper is the following theorem, which gives the global rigidity and existence of discrete conformal structures on ideally triangulated surfaces with boundary.
For convenience of description, we use the new parameters $\xi^{(1)}$ and $\xi^{(2)}$ to replace parameters $\eta$ and $\zeta$ in the mixed discrete conformal structures $\mathrm{I}$ and $\mathrm{II}$, respectively.

### v3

@@ Abstract.
In [14], we introduced discrete conformal structures on surfaces with boundary via an axiomatic framework, and provided a classification of such discrete conformal structures. The present work focuses on the rigidity and existence of these discrete conformal structures on surfaces with boundary. As a direct consequence, the results by Guo-Luo in [5] and Guo in [4], which deal with the rigidity and existence of discrete conformal structures on surfaces with boundary, are extended to a very general context.
@@ 1. Introduction
1.1. Basic definition and previous results
@@
Let $\widetilde{\mathcal{T}}=(V,\widetilde{E},\widetilde{F})$ be a triangulation of a closed surface $\widetilde{\Sigma}$, where $V$, $\widetilde{E}$, and $\widetilde{F}$ denote the sets of vertices, unoriented edges, and faces, respectively. Here $V$ is a finite subset of $\widetilde{\Sigma}$ with $|V|=N$. Let $N(V)$ be a small open regular neighborhood of the union of all vertices, chosen such that the neighborhoods around distinct vertices are disjoint. Then $\Sigma=\widetilde{\Sigma}\setminus N(V)$ is a compact surface with $N$ boundary components. The intersection $\mathcal{T}=\widetilde{\mathcal{T}}\cap\Sigma$ is referred to as an ideal triangulation of $\Sigma$. Let $E=\widetilde{E}\cap\Sigma$ and $F=\widetilde{F}\cap\Sigma$ denote the sets of unoriented ideal edges and ideal faces of $\Sigma$ with respect to $\mathcal{T}$, respectively. A boundary arc is defined as a connected component of the intersection between an ideal face and the boundary $\partial\Sigma$. For simplicity, we index the boundary components of $\Sigma$ by the set $B=\{1,2,\dots,N\}$, where each element $i\in B$ corresponds to a unique boundary component. Let $E_{+}$ denote the set of oriented ideal edges. For any two adjacent boundary components $i,j\in B$, the unoriented ideal edge connecting them is denoted by $\{ij\}\in E$, with the corresponding oriented ideal edge given by $(i,j)$. An ideal face $\{ijk\}\in F$ is adjacent to the boundary components $i,j,k\in B$. Specifically, such an ideal face is a hexagon, which corresponds to the triangle $v_{i}v_{j}v_{k}\in\widetilde{F}$ in $\widetilde{\mathcal{T}}$. The sets of real-valued functions on $B$, $E$, and $E_{+}$ are denoted by $\mathbb{R}^{N}$, $\mathbb{R}^{E}$, and $\mathbb{R}^{E_{+}}$, respectively.
@@
A edge length function associated to $\mathcal{T}$ is a vector $l\in\mathbb{R}^{E}_{>0}$ that assigns to each ideal edge $\{ij\}\in E$ a positive number $l_{ij}=l_{ji}$. For any ideal face $\{ijk\}\in F$, there exists a unique right-angled hyperbolic hexagon (up to isometry) whose three non-adjacent edges have lengths $l_{ij},l_{ik},l_{jk}$ (see [7, Theorem 3.5.14]). By gluing all such right-angled hyperbolic hexagons isometrically along the ideal edges in pairs, one can construct a hyperbolic surface with totally geodesic boundary from the ideal triangulation $\mathcal{T}$. Conversely, any ideally triangulated hyperbolic surface $(\Sigma,\mathcal{T})$ with totally geodesic boundary induces a function $l\in\mathbb{R}^{E}_{>0}$, where $l_{ij}$ is the length of the shortest geodesic connecting the boundary components $i,j\in B$. Such an edge length function $l\in\mathbb{R}^{E}_{>0}$ is referred to as a discrete hyperbolic metric on $(\Sigma,\mathcal{T})$. The length $K_{i}$ of the boundary component $i\in B$ is called the generalized combinatorial curvature of the discrete hyperbolic metric $l$ at $i$. Specifically, the generalized combinatorial curvature $K_{i}$ is defined by
(1) $K_{i}=\sum_{\{ijk\}\in F}\theta^{jk}_{i},$
@@
where the summation is taken over all right-angled hyperbolic hexagons adjacent to $i$, and the generalized angle $\theta^{jk}_{i}$ denotes the length of the boundary arc of the right-angled hyperbolic hexagon $\{ijk\}$ at $i$.
@@
Motivated by the work of Glickenstein [2] and that of Glickenstein-Thomas [3] on discrete conformal structures for triangulated closed surfaces, we introduced in [14] the following two definitions regarding partial edge lengths and discrete conformal structures on ideally triangulated surfaces with boundary.
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. An assignment of partial edge lengths is a map $d\in\mathbb{R}^{E_{+}}$ satisfying $l_{ij}=d_{ij}+d_{ji}>0$ for every edge $\{ij\}\in E$, and
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary. A discrete conformal structure $d=d(f)$ on $(\Sigma,\mathcal{T})$ is a smooth map that sends a function $f\in\mathbb{R}^{N}$ defined on the boundary components $B$ to a partial edge length function $d\in\mathbb{R}^{E_{+}}$, satisfying
@@
In general, the partial edge lengths $d\in\mathbb{R}^{E_{+}}$ may take negative values and do not satisfy the symmetry condition, i.e., $d_{ij}\neq d_{ji}$. Let $E_{ij}$ denote the hyperbolic geodesic in the hyperbolic plane $\mathbb{H}^{2}$ that extends the ideal edge $\{ij\}\in E$. The partial edge length $d_{ij}$ represents the signed distance from a point $c_{ij}\in E_{ij}$ to the boundary component $i\in B$, while $d_{ji}$ represents the signed distance from $c_{ij}$ to $j\in B$. This point $c_{ij}$ is referred to as the edge center of the ideal edge $\{ij\}$. Note that the signed distance $d_{ij}$ from $c_{ij}$ to $i$ is positive if $c_{ij}$ lies on the same side of $i$ as $j$ along the hyperbolic geodesic $E_{ij}$, and negative otherwise. Please refer to Figure 1.
Figure 1. Poincaré dual to the ideal edge $\{ij\}\in E$ with the position determined by $d_{ij}$ and $d_{ji}$ in the Klein model (dotted). The red lines represent the boundary arcs of the ideal face $\{ijk\}$, and the black triangle $v_{i}v_{j}v_{k}$ is the corresponding hyper-ideal hyperbolic triangle.
@@
We briefly elaborate on the motivation behind Definition 1.1 concerning partial edge lengths on surfaces with boundary; further details can be found in [14]. The hyperbolic geodesic passing through $c_{ij}$ and Lorentz orthogonal to $E_{ij}$ is referred to as the edge perpendicular, denoted by $P_{ij}$. This edge perpendicular $P_{ij}$ can be identified with a unique 2-dimensional time-like subspace $\widetilde{P}_{ij}$ of the Lorentzian space $\mathbb{R}^{3}$, satisfying $\widetilde{P}_{ij}\cap\mathbb{H}^{2}=P_{ij}$. Since the 2-dimensional subspaces $\widetilde{P}_{ij}$ and $\widetilde{P}_{jk}$ (associated to $P_{ij}$ and $P_{jk}$, respectively) intersect in a 1-dimensional subspace of $\mathbb{R}^{3}$, their intersection $\widetilde{P}_{ij}\cap\widetilde{P}_{jk}$ corresponds to a point $q$ in the Klein model of $\mathbb{H}^{2}$. This point $q$ is viewed as the intersection of the perpendiculars $P_{ij}$ and $P_{jk}$. Note that $q$ may be not in $\mathbb{H}^{2}$, meaning the perpendiculars $P_{ij},P_{jk}$ may not intersect within $\mathbb{H}^{2}$. Lemma 2.4 in [14] demonstrates that the perpendiculars $P_{ij}$, $P_{jk}$, and $P_{ki}$ of the ideal face $\{ijk\}$ intersect at a common point $c_{ijk}$ (termed the face center) if and only if the condition (2) is satisfied. Consequently, the condition (2) in Definition 1.1 guarantees the existence of a geometric structure on the Poincaré dual of an ideally triangulated surface with boundary. Moreover, the geometric motivation behind Definition 1.2 is discussed in detail in [14, Remark 1.3].
The main result in [14] is the following theorem, which provides explicit forms and a classification of the discrete conformal structures (as defined in Definition 1.2) on surfaces with boundary.
@@
Let $(\Sigma,\mathcal{T})$ be an ideally triangulated surface with boundary, and let $d=d(f)$ be a discrete conformal structure on $(\Sigma,\mathcal{T})$. There exist constant vectors $\alpha\in\mathbb{R}^{N}$ and $\eta\in\mathbb{R}^{E}$ satisfying $\eta_{ij}=\eta_{ji}$, such that for any right-angled hyperbolic hexagon $\{ijk\}\in F$, the value of $\frac{\sinh d_{ij}}{\sinh d_{ji}}$ can only take one of the following four forms:
$\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}},\ -\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}},\ e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})},\ -e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})},$
where $C\in\mathbb{R}^{E_{+}}$ is a constant vector satisfying $C_{ij}+C_{jk}+C_{ki}=0$ for any $\{ijk\}\in F$, and $C_{rs}+C_{sr}=0$ for any subset $\{r,s\}\subseteq\{i,j,k\}$. Furthermore,
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}>0$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then
(6) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
(A3):
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}>0$, then
(7) $\cosh l_{ij}=-\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}};$
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$ with $01+\alpha_{i}e^{2f_{i}}>$ 0 and $01+\alpha_{j}e^{2f_{j}}>$ 0, then
(8) $\cosh l_{ij}=\sqrt{(1+\alpha_{i}e^{2f_{i}})(1+\alpha_{j}e^{2f_{j}})}+\eta_{ij}e^{f_{i}+f_{j}};$
@@
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-\sqrt{\frac{1+\alpha_{i}e^{2f_{i}}}{1+\alpha_{j}e^{2f_{j}}}}<0$ with $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0, then
(B3):
if $\frac{\sinh d_{ij}}{\sinh d_{ji}}=-e^{\frac{1}{2}C_{ij}(f_{i}-f_{j})}<0$, then
(10) $\cosh l_{ij}=\cosh(f_{j}-f_{i}-C_{ij})+\eta_{ij}e^{f_{i}+f_{j}}.$
Moreover, there are exactly six types of combinations of such discrete conformal structures that can exist on the surface with boundary, namely (A1), (A2), (A3), the combination of (A1) and (B1), the combination of (A2) and (B2), and the combination of (A3) and (B3). The latter three mixed combinations are referred to as the mixed discrete conformal structure I, mixed discrete conformal structure II, and mixed discrete conformal structure III, respectively.
@@
As noted in [14, Remark 1.5], for the discrete conformal structures (A1) and (B1), the discrete conformal structure $d=d(f)$ allows a reparameterization such that $\alpha:B\to\{-1,0,1\}$, with the induced discrete hyperbolic metric $l$ remaining invariant. Thus we always assume $\alpha:B\to\{-1,0,1\}$ in (A1) and (B1) . For the discrete conformal structures (A2) and (B2), the conditions $01+\alpha_{i}e^{2f_{i}}<$ 0 and $01+\alpha_{j}e^{2f_{j}}<$ 0 imply $\alpha_{i}<0$ and $\alpha_{j}<0$. Via reparameterization, we can assume $\alpha\equiv-1$ in (A2) and (B2). For the discrete conformal structures (A3) and (B3), if $(\Sigma,\mathcal{T})$ has genus zero, a reparameterization allows us to set $C\equiv 0$. For surfaces of non-zero genus, we take $C\equiv 0$ in (A3) and (B3) as special cases.
@@
The right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ki}$ induced by (5) and (8), (6) and (9), and (7) and (10) is referred to as Type-I, Type-II, and Type-III, respectively. Let $\{r,s,t\}=\{i,j,k\}$. For a Type-I right-angled hyperbolic hexagon, if $d_{rs}<0$, then $d_{rt}<0$ by Lemma 2.16. Furthermore, from the conditions $l_{rs}=d_{rs}+d_{sr}>0$ and $l_{rt}=d_{rt}+d_{tr}>0$ in Definition 1.1, it follows that $d_{sr}>0$ and $d_{tr}>0$. For the condition (2) to hold, $d_{st}>0$ and $d_{ts}>0$ are required. Thus $l_{rs}$ and $l_{rt}$ are defined by (8), while $l_{st}$ is defined by (5). Analogous reasoning applies to Type-II and Type-III right-angled hyperbolic hexagons. Consequently, for the mixed discrete conformal structure I, all right-angled hyperbolic hexagons are either Type-I or a combination of Type-I right-angled hyperbolic hexagons and those with edge lengths induced by (5). Similar arguments also apply for the mixed discrete conformal structures II and III.
1.2. Main results
A central problem in discrete conformal geometry concerns the relationship between discrete conformal structures and their combinatorial curvatures. The main result in this paper is the following theorem, which establishes the global rigidity and existence of discrete conformal structures on ideally triangulated surfaces with boundary.